A man deposited $2000 in a savings bank paying him a certain interest. At the end of the year he withdraw $150. At the end of the second year the amount dew to him was $1947.50. What was the rate of interest?
Answer by mananth(16946)

$ 2000 in savings bank
Let x% be the interest.
..
interest obtained = 20x.
he withdrew $150
Balance in account after 1 year = 2000-150+20x = 1850+20x
...
Interest for second year = 0.01x(1850+20x)+1850+20x = 1947.50
18.50x+0.2x^2+1850+20x=1947.50
0.2x^2+38.50x=1947.50-1850
0.2x^2+38.5x=97.5
multiply by 10
2x^2+385x-975 =0
discriminant = 156025
solve by quadratic formula
x1= 2.5%
x2= -195 ignore negative value
....
m.ananth@hotmail.ca